OpenStudy (anonymous):

Inequalities help only 4 questions. Pleas. (:

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

@robtobey @zepdrix @Luigi0210 anbody who can help? @Psymon please

zepdrix (zepdrix):

Mmmmmmm ya your first answer looks good.

zepdrix (zepdrix):

Hmm not so much on the second one D: Let's see what's going on.

zepdrix (zepdrix):

\[\Large 4p+1\gt-11\]subtracting 1,\[\Large 4p\gt-12\]Dividing 4,\[\Large p\gt-3\]

zepdrix (zepdrix):

p is GREATER than -3. Did you accidently flip the inequality sign maybe? :o

OpenStudy (anonymous):

Here are my last 2.

OpenStudy (anonymous):

I dont think i did them right because this is confusing for me but if you can please explain what i did wrong and the right answer.

zepdrix (zepdrix):

\[\Large |2x+1|\le 5\]Dropping the bars gives us,\[\Large -5\le2x+1\le5\]Then solving for x gives us,\[\Large -3\le x\le 2\] Looks like you got the 3rd one correct, lol :) were you just guessing? :3

OpenStudy (anonymous):

No, i worked them out the way i was taught to but its confusing for me. I have never been good at math. so i got 3rd correct what about the 4th?

zepdrix (zepdrix):

Hmm these absolutes always confuse me also... -_- I think this is how we do it...\[\Large |2x+4|>14\]Dividing each side by 2,\[\Large |x+2|>7\]Dropping the bars let's us write it like ummmm...\[\Large (x+2)>7, \qquad\qquad (x+2)<-7\]

zepdrix (zepdrix):

Solving for x gives usssss... Yah I think you got the right answer there also! :O

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Haha yeah so freaking confusing but thanks for the help (: